OpenStudy (anonymous):

if two dice are rolled what is the probability of getting sum of 6 given that atleast one dice shows 2?

OpenStudy (cwtan):

Probability the dice shows 2=1/6 If at least one dice show two and getting the sum 6 then it will be either 4,2 or 2,4 so the answer is 2(1/6)(1/6)

OpenStudy (anonymous):

didnt get it two dice are rolled.

OpenStudy (cwtan):

now the condition: 1. Sum=6 2. At least one must be 2 so there is two possibility: 1. we get 2 first and then get 4 2.we get 4 first and then get 2 Probability of 1.: (1/6)(1/6)=1/36 Probability of 2.: (1/6)(1/6)=1/36 Total probability= 1/36+1/36=1/18

OpenStudy (cwrw238):

i'm not too good at probability but isn't this conditional probability? you are given that 1 dice is 2 so should not the 1/6 's be added?

OpenStudy (anonymous):

@cwrw238 you are right.